OpenStudy (anonymous):

3+5(x-2)=12-(x+5)+10 . Help .?

OpenStudy (shamil98):

Distribute the numbers first. 3 + 5(x-2) = 12 - (x+5) + 10 3 + 5x - 10 = 12 - x - 5 + 10 combine the like terms. 5x - 7 = -x + 17 get x on one side. 6x - 7 = 17 add 7 to both sides 6x = 24 divide by 6. x = 4

OpenStudy (anonymous):

Thanks :) Big help.

OpenStudy (anonymous):

3+5(x-2)=12-(x+5)+10 3+5x-10=12-x+5 5x+x=12+5-3+10 6x=24 6x/6=24/6 x=4

OpenStudy (anonymous):

Thanks @khadeeja

OpenStudy (anonymous):

no problem :)